OpenStudy (anonymous):

Solve: −5(x − 2) = 15 1 −3.4 −5 −1

OpenStudy (abb0t):

-5x+10 = 15 -5x = 15-10 = 5 x = -1

OpenStudy (anonymous):

thank you!